# Present address: ]Savannah River National Laboratory, Savannah River Site, SC, 29803 ]www.amberlauercoles.com

### v1

@@ I Introduction
Type I X-Ray Burst (XRB) nucleosynthesis occurs when H- and He-rich material is accreted onto the surface of a neutron star. This happens when the neutron star is in a low-mass ( M⊙) X-ray binary system with a hydrogen- and/or helium-rich donor, such as a main sequence star. The surface material then seeds the resulting nuclear reactions, which first consist of nuclear burning that occurs on the surface during accretion through the Hot Carbon-Nitrogen-Oxygen (HCNO) cycles, and then the thermonuclear runaway during the XRB. This runaway occurs when surface thermodynamic conditions reach extreme levels related to a critical accretion rate, leading to a thin-shell instability Hansen and van Horn (1975). In this mildly degenerate shell, thermal energy from nucleosynthesis is not dissipated, and instead feeds back into the rising thermodynamics, resulting in a shell-flash. The triple-alpha process drives this flash, and the energy released allows alternate reactions to bypass the time-limited steps in the HCNO cycle until finally a breakout reaction routes into the runaway thermonuclear nucleosynthesis known as the XRB. The burst is identified by the primary observable, a release of photons in the X-ray spectrum, known as the “light-curve” characterized by a rapid increase in flux, followed by an exponential decay. Type I XRB’s can last s, reach peak temperatures of GK, recur over periods of hours to days, and release ergs of energy.
The thermonuclear runaway is fueled during the rise time by the process, a series of and reactions. During the exponential decay of the burst the (rapid proton-capture) process occurs, via a series of reactions and decays ending around A Schatz and Rehm (2006). Several of the ) reactions in the former process occur on so-called waiting-point nuclei. In these even-even, nuclei the radiative proton captures and the corresponding photo-disintegration are in a equilibrium, due to their low -values. Without an alternate pathway, nucleosynthesis may stall due to the -decay half-life (1 s), which is comparable to the timescale of the burst. An reaction, once the temperature is sufficiently high, offers an alternative route for the nucleosynthesis, thus restarting the burst. Such a stalling and restarting of the nucleosynthesis may be responsible for observed double-peaked XRBs (e.g. Pennix et al. (1989)). There are four common candidates for waiting point nuclei in the process, 22Mg, 26Si, 30S, and 34Ar Fisker et al. (2004), and an additional possible candidate, 38Ca, identified in O’Brien et al. (2009). In standard, single-peaked bursts, the waiting point nuclei would still be expected to impact burst features.
@@
Due to low cross sections of the reaction and the low intensities of available 34Ar beams, the 34Ar()37K reaction has only been studied directly once by Browne et al. Browne et al. (2023), though at energies well above the astrophysically relevant energy regime. It has yet to be incorporated into rate formulations in the standard REACLIB library. As such, the theoretical rate calculation found in REACLIB Cyburt et al. (2010) is the standard value used in many stellar models. However, the underlying Hauser-Feshbach formalism used in cases where experimental values are not available, may be inappropriate if the reaction rate is dominated by a small number of resonances in the 38Ca compound nucleus Rauscher and Thielemann (2000).
A previous indirect measurement at iThemba Laboratories studied the 40CaCa reaction using the K spectrograph Long et al. (2017). This measurement confirmed 14 previously observed levels and detected more than 30 new levels between the ground state and E MeV, with of these within the Gamow widow for GK. This result agrees with measurements of the mirror nucleus, 38Ar, which contains approximately the same number of -wave resonances within this region Blackmon et al. (2014). These results were used to calculate a new 34Ar()37K reaction rate, which is lower by up to two orders of magnitude in the Gamow window when compared with TALYS V1.8 and NON-SMOKERweb/REACLIB calculations. This discrepancy is however not seen in the recent study by Browne et al. Browne et al. (2023), which used a Si detector array with the Jet Experiments in Nuclear Structure and Astrophysics (JENSA) gas jet target Chipps et al. (2014) for a direct measurement using an 34Ar beam from ReA3. Two measurements were taken at and 5.9 MeV and the resulting cross sections are consistent with the Hauser-Feschbach-calculated value of this rate. However, that study was not able to disentangle contributions from () reactions on the beam contaminants and the data obtained was above the astrophysically relevent energy regime of MeV.
Here we report on a measurement of proton scattering on 37K (Sec. II) to study properties of states in the compound nucleus, 38Ca, important for estimating the 34ArK reaction rate. The excitation spectrum resulting from this experiment and previously known resonance energies were analyzed in an -matrix calculation to constrain spin-parity values and partial widths of these resonances and to identify additional resonances previously unobserved (Sec. III). These quantities, together with theoretical and proton widths, were used to calculate the reaction rate via the narrow-resonance formalism (Sec. IV), which is compared to other existing rate calculations. It was then input into models of an XRB in the stellar evolution software MESA. The results of these models are then compared with a baseline model and previous models in Sec. V.
@@ II Experimental Design
The beam impinged on a 2.7-mg/cm2 CH2 target to produce the desired proton-scattering events. Protons emitted at were detected by Si detector telescopes arranged in quadrants as shown in Fig. 1. Two opposite quadrants consisted of a stack of two 1-mm thick Si detectors (both Micron Semiconductor model QQQ3). The other two quadrants comprised a stack of three Si detectors (Micron Semiconductor models QQQ5) with a 0.1-mm thick detector backed by two additional Si detectors of 1-mm thickness each. However, the thresholds in the ASIC electronics for detection of the low energies deposited by the protons in the 0.1-mm-thick detectors was found to be unreliable. Thus, in the subsequent analysis, data is only presented from the QQQ3 detectors. A self-consistent energy calibration was performed by combining the dominant 241Am characteristic decays ( MeV) and signals from an electronic pulser. The heavy ions (beam, contaminants, and heavy recoil products) were forward focused and passed through the center of the silicon telescopes to enter the Ionization Chamber (IC) Lai et al. (2018), which provided both position and identification by relative energy loss. The different elemental components of the beam were clearly resolved in the ionization chamber when a target was not in place, allowing the beam composition to be periodically monitored and optimized. Contaminants in the beam (13 of the beam) were nearly equal parts 37Cl and 37Ar. With the CH2 target in place, the resolution of the IC was not sufficient to discriminate scattering from the elemental components on an event-by-event basis.
@@ III Analysis
The segmented QQQ3 silicon detectors (16 rings and 16 sectors) measured the reaction angle for each event. Kinematic reconstruction converted the measured proton energy and angle of the elastic scattering events from the lab frame into center-of-mass, combining data over the entire detector into a single excitation function corresponding to an average detector lab angle of 22.2∘ (center-of-mass angle of 135.7∘). An absolute energy calibration was applied to the excitation function using proton-scattering data taken with an 40Ar stable beam and the same CH2 target. This proton-scattering spectrum is well characterized by a previous study Barnard and Kim (1961) with two distinctive resonant features that were aligned with the same features in the 40Ar stable beam spectrum to provide a two-point calibration. The final, calibrated spectrum from proton scattering on 37K, converted to excitation energy in 38Ca in the range of MeV, is shown in Fig. 2.
@@
The excitation function was analyzed using the -matrix formalism Descouvemont and Baye (2010) performed with AZURE2 Azuma et al. (2010). A channel radius of fm was used, estimated from with a suitably large fm to be outside of the range of the nuclear force for the compound nucleus. Previously observed states within the energy range shown in Fig. 2 were included with their energy values held fixed to the those reported in Long et al. (2017). A good fit to the excitation function could not be obtained using only these levels.
@@
Only natural-parity states contribute to the 34ArK reaction due to the total channel spin for the 34Ar system. While population of natural-parity states is favored in the 40CaCa reaction Long et al. (2017), unnatural parity states can be strongly populated in 37K scattering, so the need for additional states is expected. We added additional hypothetical levels to the -matrix model with energies initially set to the most prominent features in the data. This resulted in a total of 24 states being included in the analysis (13 of which are newly identified here), which is consistent with the level density found in the mirror nucleus 38Cl. With few experimental constraints on resonance properties, the resulting large size of the parameter space and likelihood for many local minima required implementing some assumptions in the analysis.
@@
We included only proton partial widths corresponding to protons populating the ground state in 37K. The lowest-energy part of the excitation function ( MeV) may have some contamination from proton inelastic scattering originating from the upstream half of the target that populates the first-excited state in 37K at 1.37 MeV. However, the observed yield of protons with energies too low to correspond to elastic scattering indicates that this contribution is likely negligible.
@@
Only the lowest orbital angular momentum value was included for any state’s particular , but both possible channel-spin combinations were included de Boer (2021). We initially fixed all the resonance energies. Initial values for partial widths were chosen that were comparable to the size of features in the excitation function or the experimental resolution from Ref. Long et al. (2017). The excitation function was fit using MINUIT2 varying partial widths constrained to be within a limited range, varying by no more than a factor of 1.7. Once a local minimum was obtained within this restricted parameter space, the value of a randomly selected state was changed, and the gradient fit repeated. The process of randomly changing a value and repeating a gradient fit was iterated, starting each cycle with the best previous fit, until no improvement was gained.
@@ IV Reaction rate
The reaction rate is calculated using the narrow resonance approximation from the definition provided in Iliadis (2008), with the total reaction rate expressed as a sum over individual resonances described by the Breit-Wigner formalism:
@@
Here, and are the are partial widths of the entrance () and exit (p) channels, respectively; is the total width of the resonance ( in this case), and , and are the spins of the resonance and the reactants ( and 34Ar in this case with ), respectively. While the -matrix formalism would be more appropriate given the level density in this case, the narrow-resonance approach of Long et al. (2017) is followed to allow a more direct comparison of the rates and impact.
The partial width is related to the single-particle width, , thusly: =C, where C is the spectroscopic factor. The same constant spectroscopic factor of 0.01 used in Ref. Long et al. (2017) is used here for the partial widths for consistency. In addition to this constant factor, a distribution of spectroscopic factors set to mimic the existence of -cluster states was used in a separate analysis by Long et al. (2017); however, this structure has not been observed and thus is not explored in this work. Proton partial widths from the R-matrix analysis discussed above and estimated gamma partial widths were used to verify that , implying resonance strengths can be accurately estimated using .
The total reaction rate calculated here includes the new resonance information as well as additional levels reported in Ref. Long et al. (2017); Long (2016), as the latter includes levels outside the experimental region explored here that still contribute to the overall reaction rate. The reaction rate was calculated times using a Monte-Carlo sampling of spin assignments for each resonance. The distribution of spins was taken from Long et al. (see Fig. 3 in Long et al. (2017)).
Statistical analysis provides the recommended median rate, along with the 1st and 3rd quartile. The uncertainties are dominated by the uncertain spin-parities and alpha-partial widths of the states of interest. The resulting rate is presented in Fig. 3 along with the rates from the previous work by Long et al. (2017); Long (2016), as well as REACLIB Cyburt et al. (2016). There is good agreement between the astrophysical reaction rate calculated in this work and Long et al. Long et al. (2017), varying by an average factor of .915, while both vary from the REACLIB Hauser Feschbach-based rate by approximately a factor of 45. This, however, disagrees with the conclusions of Browne et al. Browne et al. (2023), which state that the Hauser-Feschbach estimate is likely sufficient for this reaction.
@@ V Impact on Stellar Models
The ultimate impact of this new reaction rate was tested using a Type-I X-ray Burst model adapted from Paxton et al. Paxton et al. (2015), Modules for Experiments with Stellar Astrophysics (MESA), version 10108. MESA is a multi-zone, 1D radial stellar evolution code, with adaptive mesh in time and space that includes relevant physics such as opacities, hydrodynamics, mixing length theory, atmospheric boundary conditions, and equation of state. The nuclear network differential equations are coupled to those that model the physical aspects so that the energy generation provides feedback into the thermodynamics and vice versa, and includes mixing between zones. The model is designed to simulate conditions of a hydrogen-dominant burst on GS 1826-24, colloquially known as the “clocked burster” due to its extreme regularity Galloway et al. (2004), with an initial neutron-star mass and radius of 1.4 M⊙ and 11.2 km, respectively. This model assumes solar abundances for accretion (1H , 4He , and from Grevesse and Sauval (1998)) at a rate of Myr. The nuclear network is similar to the classic 304-species XRB network identified by Fisker et al. (2008), favoring proton-rich isotopes up to 107Te, but with the neutron-degenerate core modelled as the heavy element 138Ba, which is technically included in the nuclear network but is far above the next-greatest isotope and so unreachable by any standard capture reaction. The result is that the core is essentially inert. Additional astrophysical details of the model can be found in Ref. Paxton et al. (2015).
@@
Here we report the results of 11 variations of the base model to test a sample of the range of known burst conditions with different 34Ar()37K reaction rates. For five models there was no alteration save the 34Ar()37K reaction rate. This set included the standard REACLIB rate, REACLIB and , the recommended rate from this work, and REACLIB (i.e. approximately the average ratio of this work to REACLIB). These factors were chosen to compare with similar published stellar modelling studies. The standard REACLIB rate is from the 2010 compilation of theoretical Hauser-Feschbach rate calculations Cyburt et al. (2010).
@@
Three additional models increased the He accretion percentage to to mimic a He-rich companion and provide conditions where an -induced reaction might be especially impactful. Three more models tested a slowed accretion (M⊙/yr) compared to the “clocked burster” model. These six models all used the 34Ar()37K REACLIB rate as the baseline with the same variations by factors of and .
Finally, an additional three models changed the metallicity of the accreted material to a relatively high value of 0.19, about ten times Solar metallicity. This was done for comparison to the astrophysical scenario used in Parikh et al. Parikh et al. (2008), where the high-Z starting composition model resulted in the largest impact from varying the 34Ar()37K rate, altering the resulting abundance of 34S. In these models the starting metallicity was of a similar value as Parikh et al., albeit slightly lower in order to produce a model that was numerically stable. Indeed, this group of models only exhibited one burst, after which the model was quiescent for an extended time and then unable to find suitable solutions to the differential equations such that evolution did not proceed and thus the results are not discussed here. It should be noted that Fisker et al. treated this high- scenario in the initial abundance, rather than the accretion, a scenario which is outside the scope of this work. It should also be noted that the single burst produced in this model was the only output light curve that exhibited a double peak.
ch is why I made up my own.
@@
The models investigated here were compared via the following metrics: burst period (absolute and differential), integrated luminosity (definition given below), and maximum luminosity (absolute and differential). A summary of the results of the stellar modelling is given in Table 2. The number of bursts is also listed, and gives a general sense of the numerical stability of the model, with more bursts corresponding to higher stability. The full set of stable burst models as well as the folded, averaged burst light curves, which are used to calculate integrated luminosity metric, are reflected in Figs. 4, 5, and 6. Final abundances were also tracked for this study; however, there were no significant differences in the final abundances produced by the bursts with the variation of the 34Ar()37K reaction rate.
@@
The three different sets of models result in differences in the burst duration from each other, but as this was not caused by variations in the 34Ar()37K rate, it is not relevant to this work. The only significant change in burst duration due to this rate variation was in the model using slower accretion (see Fig. 5) with the 34Ar()37K rate set at the standard REACLIB rate, which was 2689 s, or 9, longer. The same model with the rate varied by showed a slightly smaller variation, while other models varied by . The absolute maximum luminosity showed more significant variations in the models tested, with the largest variation of 260% again in the slower accretion model with the reaction rate set at 0 the REACLIB rate, followed by the slow accretion model with reaction rate at (120%), and the high He-accretion models with the rate at (22%) and (9%).
@@
where is a baseline model with accepted rates from REACLIB Cyburt et al. (2010). This quantity was similar for all of the “clocked burster” variations. Though these models had similar burst periods and maximum luminosities, variations in burst and quiescent period shape and values could explain these non-zero values (i.e. the differences from the baseline model). The high He-accretion model using the standard reaction rate had the largest value. As this model also had a larger difference in maximum luminosity, this is easier to explain. Both slow accretion models had modest values for this metric.
| Model Identifier | no. bursts | burst period(s) | burst diff% | erg | erg/s | diff |

### v2

# Present address: ]Savannah River National Laboratory, Savannah River Site, SC 29803, USA ]www.amberlauercoles.com

@@ I Introduction
Type I X-Ray Burst (XRB) nucleosynthesis occurs when H- and He-rich material is accreted onto the surface of a neutron star. This happens when the neutron star is in a low-mass X-ray binary system with a hydrogen- and/or helium-rich donor, such as a main sequence star. The surface material then seeds the resulting nuclear reactions, which first consist of nuclear burning that occurs on the surface during accretion through the Hot Carbon-Nitrogen-Oxygen (HCNO) cycles, and then the thermonuclear runaway during the XRB. This runaway occurs when the temperature-dependent triple- reaction rate increases, significantly contributing energy to the system and leading to a thin-shell instability Hansen and van Horn (1975). In this mildly degenerate shell, thermal energy from nucleosynthesis is not dissipated, and instead feeds back into the rising temperature, resulting in a shell-flash. The rising temperature allows additional nuclear reactions that lead to breakout of the hCNO cycles and subsequent nucleosynthesis via the and processes giving rise to the X-ray burst. The burst is identified by the primary observable, a release of photons in the X-ray spectrum, known as the “light-curve”, characterized by a rapid increase in flux, followed by an exponential decay. Type I XRBs can last s, reach peak temperatures of GK in the nuclear burning layer, recur over periods of hours to days, and release approximately ergs of energy.
The XRB is fueled during the initial rise time by the process, a series of and reactions, followed by the exponential decay of the burst driven by the (rapid proton-capture) process, a series of reactions and decays ending around A Schatz and Rehm (2006). Several of the ) reactions in the former process occur on so-called waiting-point nuclei. In these even-even, nuclei the radiative proton captures and the corresponding photo-disintegration are in a equilibrium, due to their low -values. Without an alternate pathway, nucleosynthesis may stall due to the relatively long -decay half-life (1 s), which is comparable to the timescale of the burst. An reaction, once the temperature is sufficiently high, offers an alternative route for the nucleosynthesis, thus restarting the burst. Such a stalling and restarting of the nucleosynthesis may be responsible for observed double-peaked XRB light curves (e.g. Pennix et al. (1989)). There are four common candidates for the dominant waiting point nuclei in the process, 22Mg, 26Si, 30S, and 34Ar Fisker et al. (2004), and an additional possible candidate, 38Ca, identified in Ref. O’Brien et al. (2009). In standard, single-peaked bursts, the waiting point nuclei have also been shown to impact burst features.
@@
Due to low cross sections of the reaction and the low intensities of available 34Ar beams, the 34Ar()37K reaction has only been studied directly once by Browne et al. Browne et al. (2023), though at energies well above the astrophysically relevant energy regime. It has yet to be incorporated into rate formulations in the standard REACLIB library Cyburt et al. (2010). As such, the theoretical rate calculated with the NON-SMOKER statistical model Rauscher (1999) and found in REACLIB is the standard value used in many stellar models. However, the underlying Hauser-Feshbach formalism used in cases where experimental values are not available, may be inappropriate if the reaction rate is dominated by a small number of resonances in the 38Ca compound nucleus Rauscher and Thielemann (2000).
A previous indirect measurement at iThemba Laboratories studied the 40CaCa reaction using the K spectrograph Long et al. (2017). This measurement confirmed 14 previously observed levels and detected more than 30 new levels between the ground state and E MeV, with of these within the Gamow widow for GK. This result agrees with measurements of the mirror nucleus, 38Ar, which contains approximately the same number of -wave resonances within this region Blackmon et al. (2014). These results were used to calculate a new 34Ar()37K reaction rate, which is lower by up to two orders of magnitude in the Gamow window when compared with TALYS-1.8 Koning and Rochman (2012) and NON-SMOKERweb/REACLIB calculations. However, this discrepancy was not seen in the recent study by Browne et al. Browne et al. (2023), which used a Si detector array with the Jet Experiments in Nuclear Structure and Astrophysics (JENSA) gas jet target Chipps et al. (2014) for a direct measurement using an 34Ar beam from ReA3. Two measurements were taken at and 5.9 MeV and the resulting cross sections are consistent with the Hauser-Feschbach-calculated value of this rate. However, that study was not able to disentangle contributions from () reactions on the beam contaminants and the data obtained was above the astrophysically relevent energy regime of MeV.
Here we report on a measurement of proton scattering on 37K (Sec. II) to study properties of states in the compound nucleus, 38Ca, important for estimating the 34ArK reaction rate. The excitation spectrum resulting from this experiment and previously known resonance energies were analyzed in an -matrix calculation to constrain spin-parity values and partial widths of these resonances and to identify additional resonances previously unobserved (Sec. III). These quantities, together with theoretical and proton widths, were used to calculate the reaction rate via the narrow-resonance formalism (Sec. IV), which is compared to other existing rate calculations. It was then input into models of an XRB in the stellar evolution software MESA (Modules for Experiments in Stellar Astrophysics). The results of these models are then compared with a baseline model and previous models in Sec. V.
@@ II Experimental Design
The beam impinged on a 2.7-mg/cm2 CH2 target to produce the desired proton-scattering events. Protons emitted at were detected by Si detector telescopes arranged in quadrants as shown in Fig. 1. Two opposite quadrants consisted of a stack of two 1-mm thick Si detectors (both Micron Semiconductor model QQQ3). The other two quadrants comprised a stack of three Si detectors (Micron Semiconductor models QQQ5) with a 0.1-mm thick detector backed by two additional Si detectors of 1-mm thickness each. However, the thresholds in the ASIC electronics for detection of the low energies deposited by the protons in the 0.1-mm-thick detectors were found to be unreliable. Thus, in the subsequent analysis, data is only presented from the QQQ3 detectors. A self-consistent energy calibration was performed by combining the dominant 241Am characteristic decays ( MeV) and signals from an electronic pulser. The heavy ions (beam, contaminants, and heavy recoil products) were forward focused and passed through the center of the silicon telescopes to enter the Ionization Chamber (IC) Lai et al. (2018), which provided both position and identification by relative energy loss. The different elemental components of the beam were clearly resolved in the ionization chamber when a target was not in place, allowing the beam composition and intensity to be periodically monitored and optimized. Contaminants in the beam (13 of the beam) were nearly equal parts 37Cl and 37Ar. With the CH2 target in place, the resolution of the IC was not sufficient to discriminate scattering from these elemental components on an event-by-event basis; however, the scattering cross section for these contaminants is not expected to vary significantly compared to the 37K elastic scattering cross section.
@@ III Analysis
The segmented QQQ3 silicon detectors (16 rings and 16 sectors) measured the reaction angle for each event. Kinematic reconstruction converted the measured proton energy and angle of the elastic scattering events from the lab frame into center-of-mass, combining data over the entire detector into a single excitation function corresponding to the average cross section over the angular range of . An absolute energy calibration was applied to the excitation function using proton-scattering data taken with an 40Ar stable beam and the same CH2 target. This proton-scattering spectrum is well characterized by a previous study Barnard and Kim (1961) with two distinctive resonant features that were aligned with the same features in the 40Ar stable beam spectrum to provide a two-point calibration. The final, calibrated spectrum from proton scattering on 37K, converted to excitation energy in 38Ca in the range of MeV (E MeV), is shown in Fig. 2.
@@
The excitation function was analyzed using the -matrix formalism Descouvemont and Baye (2010) performed with AZURE2 Azuma et al. (2010). A channel radius of fm was used, estimated from with a suitably large fm to be outside of the range of the nuclear force for the compound nucleus. The excitation function was fit with the calculated differential cross section at a single (average) angle of to reduce computational complexity, but this is a good approximation given the limited angular range and modest angular dependence. Previously observed states within the energy range shown in Fig. 2 were included with their energy values held fixed to those reported in Long et al. (2017). The success of an AZURE2 fit was judged by the , trying to minimize this value, while reflecting the impact of the resonance energy, partial width, and spin-parity values.
A good fit to the excitation function could not be obtained using only the levels identified in Long et al. The fit using only those levels resulted in a of 2.263 and regions of the fit that were not representative of the corresponding features of the experimentally determined cross section.
@@
Only natural-parity states contribute to the 34ArK reaction due to the total channel spin for the 34Ar system. While population of natural-parity states is favored in the 40CaCa reaction Long et al. (2017), unnatural parity states can be strongly populated in 37K scattering, so the need for additional states is expected. We added additional hypothetical levels to the -matrix model with energies initially set to the most prominent features in the data. This resulted in a total of 24 states being included in the analysis (13 of which are newly identified here), which is consistent with the level density found in the mirror nucleus 38Cl.
@@
With few experimental constraints on resonance properties, the resulting large size of the parameter space and likelihood for many local minima required implementing some assumptions in the analysis. We included only proton partial widths corresponding to protons populating the ground state in 37K. The low-energy half of the excitation function ( MeV) may have some contamination from proton inelastic scattering originating from the upstream half of the target that populates the first-excited state in 37K at 1.37 MeV. However, the observed yield of protons with energies too low to correspond to elastic scattering indicates that this contribution is likely negligible.
@@
Only the lowest orbital angular momentum value was included for any state’s particular , but both possible channel-spin combinations were included de Boer (2021). We initially fixed all the resonance energies. Initial values for proton partial widths were chosen that were comparable to the width of features in the excitation function or the experimental resolution from Ref. Long et al. (2017), which sets an upper limit on the proton widths for those observed states. The excitation function was fit using MINUIT2 varying partial widths constrained to be within a limited range, varying by no more than a factor of 1.7. Once a local minimum was obtained as demonstrated by the within this restricted parameter space, the value of a randomly selected state was changed, and the gradient fit repeated. The process of randomly changing a value and repeating a gradient fit was iterated, starting each cycle with the best previous fit, until no improvement was gained.
@@ IV Reaction rate
The reaction rate is calculated using the narrow resonance approximation from the definition provided in Iliadis (2008), with the total reaction rate (in cm3mol-1s-1) expressed as a sum over individual resonances described by the Breit-Wigner formalism:
@@
Here, and are the are partial widths of the entrance () and exit (p) channels, respectively; is the total width of the resonance ( in this case), and , , and are the spins of the resonance and the reactants ( and 34Ar in this case with ), respectively. The partial width is related to the single-particle width, , thusly: =C, where C is the spectroscopic factor. While an -matrix formalism would be more appropriate given the level density, the narrow-resonance approach of Long et al. (2017) is followed using the same constant to allow direct comparison to the previous work. Proton partial widths from the -matrix analysis discussed above and estimated gamma partial widths were used to verify that , implying resonance strengths can be accurately estimated using .
The total reaction rate includes the new resonance information from this work as well as additional levels reported in Ref. Long et al. (2017); Long (2016), as the latter includes levels outside the experimental region explored here that still contribute to the overall reaction rate. The reaction rate was calculated times using a Monte-Carlo sampling of natural-parity spin assignments for each resonance taken from Long et al. (see Fig. 3 in Long et al. (2017)) outside of our experimentally measured energy region. For states between E MeV, the lowest natural parity assignment from Table 1 was used or the state was not included if the only available spin assignment(s) was of unnatural parity. The median rate, which is the sum of these contributions, with the 1st and 3rd quartile of rates are presented in Fig. 3 along with the rates from the previous work by Long et al. (2017); Long (2016). We also include the rate from REACLIB Cyburt et al. (2010) based on NON-SMOKER Rauscher (1999), which is nearly the same as the statistical model rate from TALYS-1.8 Koning and Rochman (2012) using the default model parameters.
Our rate is higher than the rate of Long et al. Long et al. (2017) for GK due to the inclusion of additional levels but generally consistent at higher temperatures. However, compared to the statistical model rates, our rate is lower by a factor of . Given that the density of states in the region is close to that of the mirror nucleus, it is unlikely that a large number of states are missing. The simple assumption of may lead to an underestimate of the rate, particularly as -cluster states could result in significantly higher , though likely less than the observed discrepancy with the statistical model predictions Long et al. (2017). Direct measurement of the 34Ar()37K total cross section at higher energies found good agreement with statistical model predictions in the regime where the reaction proceeds to many excited final states. However, population of the ground state in 37K was found to be much weaker than model predictions Browne et al. (2023). Population of the ground state is expected to dominate at the energies covered in this measurement (and much of the Gamow window), and indeed is the only state that can be populated in the lower part of the Gamow window. Thus, the reliability of the statistical model in this lower energy regime is less clear, and uncertainties in the reaction rate are still substantial.
@@ V Impact on Stellar Models
Given the uncertainty in the rate, we re-examined the impact of the 34Ar()37K reaction rate using a Type-I X-ray Burst model adapted from Paxton et al. Paxton et al. (2015), Modules for Experiments with Stellar Astrophysics (MESA), version 10108. MESA is a multi-zone, 1D radial stellar evolution code, with adaptive mesh in time and space that includes relevant physics such as opacities, hydrodynamics, mixing length theory, atmospheric boundary conditions, and equation of state. The nuclear network differential equations are coupled to those that model the physical aspects so that the energy generation provides feedback into the thermodynamics and vice versa, and includes mixing between zones. The model is designed to simulate conditions of a hydrogen-dominant burst on GS 1826-24, colloquially known as the “clocked burster” due to its extreme regularity Galloway et al. (2004), with an initial neutron-star mass and radius of 1.4 M⊙ and 11.2 km, respectively. This model assumes solar abundances for accretion (1H , 4He , and from Grevesse and Sauval (1998)) at a rate of Myr. The nuclear network is similar to the classic 304-species XRB network identified by Fisker et al. (2008), favoring proton-rich isotopes up to 107Te; however, the base layer down to the degenerate core is modeled as the heavy element 138Ba, which is technically included in the nuclear network but is far above the next-greatest isotope and so unreachable by any standard capture reaction, making the core inert. Additional astrophysical details of the model can be found in Ref. Paxton et al. (2015).
@@
Here we report the results of 11 variations of the base model to test a sample of the range of known burst conditions with different 34Ar()37K reaction rates. For five models there was no alteration save the 34Ar()37K reaction rate. This set included the standard REACLIB rate, REACLIB and , the recommended rate from this work, and REACLIB (i.e. approximately the average ratio of this work to REACLIB). These factors were chosen to compare with similar published stellar modeling studies. The standard REACLIB rate is from the 2010 compilation of theoretical Hauser-Feschbach rate calculations Cyburt et al. (2010). The MESA models were run for 3000 timesteps and all analysis values were calculated over these, except for abundances which were calculated after quiescence at the 10th burst.
@@
Three additional models increased the He accretion percentage to to mimic a He-rich companion and provide conditions where -induced reactions might be especially impactful. Finally, three more models tested a slowed accretion (M⊙/yr) compared to the “clocked burster” model. These six models all used the 34Ar()37K REACLIB rate as the baseline with the same variations by factors of and .
@@
The models investigated here were compared via the following metrics: burst period (absolute and differential), integrated luminosity difference (definition given below), and maximum luminosity (absolute and differential). A summary of the results of the stellar modelling is given in Table 2. The number of bursts is also listed, and gives a general sense of the numerical stability of the model, with more bursts corresponding to higher stability. The full set of stable burst models as well as the folded, averaged burst light curves, which are used to calculate integrated luminosity metric, are reflected in Figs. 4, 5, and 6.
@@
The three different sets of models result in differences in the burst frequency from each other, but as this was not caused by variations in the 34Ar()37K rate, it is not relevant to this work. The only significant change in burst frequency due to this rate variation was in the model using slower accretion (see Fig. 5) with the 34Ar()37K rate set at the standard REACLIB rate, which was 2689 s (9) longer. The same model with the rate varied by showed a slightly smaller variation, while other models varied by . The absolute maximum luminosity showed more significant variations in the models tested, with the largest variation of 260% again in the slower accretion model with the reaction rate set at the REACLIB rate, followed by the slow accretion model with reaction rate at (120%), and the high He-accretion models with the rate at (22%) and (9%).
@@
where is a baseline model with accepted rates from REACLIB Cyburt et al. (2010). This quantity was similar for all of the “clocked burster” variations, which differed by less than 13% between models. Though these models had similar burst periods and maximum luminosities, variations in burst and quiescent period shape and values could explain these non-zero values (i.e. the differences from the baseline model). The high He-accretion model using the standard reaction rate had the largest value, likely due to the larger difference in maximum luminosity compared with the Clocked-Burster series, along with more variation throughout the burst tail. Both slow accretion models had modest values for this metric (Table 2).
| Model Identifier | no. bursts | burst period (s) | burst diff (%) | erg) | erg/s) | diff () |